Triangle ABC is a right triangle. Side AC has a length of 8M and side BC has a length of 4M. Find the length of side AB.
Answer by Cromlix(4381)

I am assuming that your right angle is at B
AB^2 + BC^2 = AC^2
AB^2 = AC^2 - BC^2
AB^2 = (8)^2 - (4)^2
AB^2 = 64 - 16
AB = square root of 48
AB = 6.9 ( 1 decimal place)